OpenStudy (anonymous):

3x + 14x - 5 =0 solve using the quadratic formula

OpenStudy (jadeishere):

is 3x squared?

OpenStudy (jadeishere):

we need one term squared (a) one variable not squared (b) and one coefficient (c)

OpenStudy (anonymous):

yeah

OpenStudy (jadeishere):

so it is squared? brilliant

OpenStudy (anonymous):

yeah typo

OpenStudy (jadeishere):

A = 3 B = -14 C = -5 Accordingly, B2 - 4AC = 196 - (-60) = 256 14 ± √ 256 x = —————— 6 The prime factorization of 256 is 2•2•2•2•2•2•2•2 √ 256 = √ 2•2•2•2•2•2•2•2 =2•2•2•2•√ 1 = ± 16 • √ 1 = ± 16 x = ( 14 ± 16) / 6 x =(14+√256)/6=(7+8)/3= 5.000 and x =(14-√256)/6=(7-8)/3= -0.333 x = -1/3 and x= 5

OpenStudy (anonymous):

you are a goddess! I thank you!

OpenStudy (jadeishere):

I wouldn't go that far xD Anytime :)